A rectangular field has to be fenced on three sides leaving a side of 20 feet uncovered. If the area of the field is 680 sq. feet, how many feet of fencing will be required?



A. 95



B. 92



C. 88



D. 82


Grade:upto college level

4 Answers

Hrishant Goswami
10 years ago

Answer : Option C

Explanation :

Given that area of the field = 680 sq. feet
=> lb = 680 sq. feet

Length(l) = 20 feet

=> 20 × b = 680

B=680/20=34 feet

Required length of the fencing = l + 2b = 20 + (2 × 34) = 88 feet

Aneesh
10 years ago

It is Option C. - 88

hari
10 years ago

opoction  _c       88

PERUMAL
10 years ago
option c 88